On January 1,2021,Katty Company issued$800,000 of 8%,10-year bonds for 97.Patty retired all of these bonds on January 1,2022,at 102.If Katty uses the straight-line amortization,how
much loss should be recognized on this bond retirement?
Answer #1

Par value of bonds = $800,000

Issue price =97

Cash received from issuance of bonds = Par value of bonds x Issue price

= 800,000 x 97%

= $776,000

Discount on issue of bonds = Par value of bonds- Cash received from issuance of bonds

= 800,000-776,000

= $24,000

Annual amortization of bond discount = Discount on issue of bonds / Bond life

= 24,000/10

= $2,400

Unamortized bond discount on January 1, 2022 = Discount on issue of bonds - Annual amortization of bond discount

= 24,000-2,400

= $21,600

Carrying value of bonds on January 1, 2022 = Par value of bonds - Unamortized bond discount on January 1, 2022

= 800,000-21,600

= $778,400

Cash paid to retire bonds = Par value of bonds x Call price

= 800,000 x 102%

= $816,000

Loss on bond retirement = Cash paid to retire bonds- Carrying value of bonds on January 1, 2022

= 816,000-778,400

= $37,600

Kindly comment if you need further assistance. Thanks‼!